### v1

# A Novel Approach to Industrial Defect Generation through Blended Latent Diffusion Model with Online Adaptation

Abstract
Effectively addressing the challenge of industrial Anomaly Detection (AD) necessitates an ample supply of defective samples, a constraint often hindered by their scarcity in industrial contexts. This paper introduces a novel algorithm designed to augment defective samples, thereby enhancing AD performance. The proposed method tailors the blended latent diffusion model for defect sample generation, employing a diffusion model to generate defective samples in the latent space. A feature editing process, controlled by a ”trimap” mask and text prompts, refines the generated samples. The image generation inference process is structured into three stages: a free diffusion stage, an editing diffusion stage, and an online decoder adaptation stage. This sophisticated inference strategy yields high-quality synthetic defective samples with diverse pattern variations, leading to significantly improved AD accuracies based on the augmented training set. Specifically, on the widely recognized MVTec AD dataset, the proposed method elevates the state-of-the-art (SOTA) performance of AD with augmented data by 1.5%, 1.9%, and 3.1% for AD metrics AP, IAP, and IAP90, respectively. The implementation code of this work can be found at the GitHub repository https://github.com/GrandpaXun242/AdaBLDM.git
Index Terms:
Anomaly detection, Blended latent diffusion model, Online adaption.
I Introduction
In a practical manufacturing workflow, obtaining defective samples is considerably more challenging compared to acquiring defect-free ones. Consequently, the majority of recently proposed algorithms for industrial defect inspection, such as [1, 2, 3, 4, 5, 6, 7], address the issue as an Anomaly Detection (AD) problem—an established machine learning challenge [8, 9].
In these industrial AD algorithms, the assumption is made that all anomaly-free samples, be they complete images or image patches, belong to a single distribution. Meanwhile, defective samples are identified as “outliers”. The strategic choice of adopting an Anomaly Detection approach for industrial defect inspection is evident: the absence of the need for defective samples during the training stage makes these AD-based algorithms inherently compatible with practical manufacturing scenarios.
Figure 1: Illustration of three defect generation styles. From top to bottom: conventional approaches, GAN-based algorithms, and the proposed method.
On the flip side, achieving this compatibility comes at the expense of an extremely imbalanced data distribution, a characteristic that is deemed unacceptable in the context of most discriminative algorithms. Consequently, some researchers suggest the generation of artificial defect patterns using conventional approaches [1, 3, 5] to facilitate subsequent discriminative learning processes.
More recently, state-of-the-art (SOTA) algorithms [10, 11, 12] highlight that, in addition to synthetic defects, incorporating a few real anomalous samples during the training stage can significantly enhance Anomaly Detection (AD) accuracy. This approach holds the promise of achieving even higher performance when superior simulated defective patterns are integrated into training. Pioneering works [13, 14] have already demonstrated the benefits of generating these ”lifelike” defects using more sophisticated methods. The first two rows of Fig. 1 illustrate the existing two styles of defect generation.
In this paper, we propose the integration of the cutting-edge AI generation algorithm, namely, the Diffusion Probabilistic Model (DPM) [15], into the realm of defect generation. Specifically, we leverage the Blended Latent Diffusion Model [16, 17, 18] as our foundational framework. To enhance the generation of defective samples, we implement three key modifications to the original Blended Latent Diffusion Model (BLDM). Firstly, we design a novel “defect trimap” to delineate the target object mask and the defective regions on the generated image, incorporating it as a new form of controlling information alongside the language prompts within the stable diffusion model. Secondly, to ensure the authenticity of the generated samples, a cascade “editing” stage is introduced for the BLDM, operating in both the latent and pixel space. Finally, an innovative online adaptation of the image encoder is proposed to refine the quality of the generated images. Consequently, the highly customized BLDM algorithm, referred to as AdaBLDM in this paper, exhibits the ability to generate defective samples with significantly higher image quality and pattern variations compared to those produced by state-of-the-art (SOTA) methods. The third row of Fig. 1 provides a brief illustration of the proposed BLDM-based method. In summary, the contributions of this paper are threefold, as outlined below.
  • To the best of our knowledge, it is the first time that the Diffusion Model has been adopted for the task of industrial defect generation. The advantages of the BLDM such as the generation stability, the high image quality, and the controllable content could all contribute to achieving a better defect generator compared with the SOTA methods.
  • To tailor the BLDM algorithm to the defect generation scenario, we bake BLDM in the aspects of controlling information, preheating stage, and online adaptation, respectively. This endeavor leads to the proposed algorithm, termed AdaBLDM, which achieves notably improved defective samples compared to synthetic ones generated by existing SOTA methods.
[…]
The remainder of this paper is organized as follows. In Sec. II, we discussed the related work in literature. The section III presents the proposed method in detail. Experimental results and the ablation study are addressed in Sec. IV. We conclude this paper in Sec. V.
II Related work
II-A Synthetic Defect Generation
As introduced in Sec. I, due to a scarcity of defective samples, many researchers [1, 2, 3, 5, 6] are “compelled” to generate synthetic defective samples for better AD performance. The most straightforward way to mimic defects is to “paste” anomaly pixels on the normal images as a special case of data augmentation. Baseline methods like [20], CutPaste [21], and [1] randomly cut regions from normal images and paste them to the “incorrect” places as artificial defects. Crop&Paste [22] and PRN[23] crop defect areas from genuine defective images and paste them onto defect-free images. Those algorithms, though achieve better AD performance compared with the standard one-class AD approaches, can not create new defect patterns and thus likely lead to overfitting problems. To increase the variations of anomalies, DRAEM [3], DeSTseg [6], MegSeg [5] and ReSynthDetect [24] employ extra datasets combined with Berlin noise to fabricate defects. However, in these cases, the distribution of fabricated defects differs from the real one and thus the performance gain can not be ensured.
Encouraged by the great success of AIGC (AI Generated Contents) methods including GAN (Generative Adversarial Network) algorithms [25] and Diffusion Models [26], some efforts have been made to generate more natural-looking and various simulated defects. In particular, SDGAN [27] employs two generators to switch the defective and non-defective status for sample images. SDGAN can generate a high-quality, diverse dataset of steel surface with both defective and defect-free images, thus higher classification performance is obtained. Similarly, Defect-GAN [28] can convert non-defective images into defective ones to enhance the learning process of the defect classifier. However, the above two GAN-based generation algorithms can not generate well-defined defective regions. Consequently, they could boost the performance of defect classification while can hardly lift the defect segmentation accuracy which is crucial in many real-world applications.
To the best of our knowledge, [13] and [14] are the only two exiting works that generate both the synthetic defective images as well as the corresponding pixel-level labels. In specific, the DCDGANc algorithm proposed in [13] design a novel GAN-based method to mimic real defects and the defects are fused with defect-free images via an improved Poisson blending algorithm. On the other hand, [14] introduced DFMGAN, stems from StyleGANv2 [29], to generate defect images and defect masks by using the proposed defect-aware residual blocks. Although the above two GAN-based methods can generate high-quality images with defects, they fall short in generating the defect mask strictly aligned to the generated defective pixels on the image. The misalignment stems from the disparate nature of generative methods which is a blessing in image diversity whereas it could significantly mislead the defect segmentation model in the scenario of AD. In contrast, in this paper, we propose a novel defect generation method based on Diffusion Models. We tame this AIGC algorithm to generate diverse, realistic but accurately-controlled defective image samples via a series of novel modifications.
II-B Diffusion Probabilistic Models for Image Editing
Compared with the more traditional generation methods like Generative Adversarial Networks (GANs) [30], Variational AutoEncoders (VAEs) [31] and flow-based algorithms [32], Diffusion Probabilistic Models (DPMs) [15, 33, 34, 35] enjoy a higher quality of the synthetic images and impressive stability of training. Accordingly, during the past few years, DPM-based approaches, such as Stable Diffusion (Stability AI) [36], DALL-E2 (OpenAI) [37], and Imagen (Google) [38], have achieved state-of-the-art performances in the task of content generation.
In the subdomain of image editing, DPM-based methods also play the most important role nowadays. In particular, the Latent Diffusion Model [16] conducts the image generation or editing in the latent space with a much less feature dimensionality and thus achieve much higher editing efficiency. RePaint [39] fully leverages a pre-trained DDPM [34] to perform image editing by sampling given pixels during the reverse diffusion stage. DCFace [40] employs two feature encoders to control the Diffusion Model, facilitating facial image editing. [41] proposes the ControlNet to govern existing Diffusion models to achieve a better-controlled synthesis. [17, 18] enables image editing by making modifications in either the pixel space or feature space. [42] introduces sparse text optimization to achieve better image editing.
Those methods, though achieve SOTA performance in general image editing tasks, the fuzzily governed synthesis easily fails to generate the image regions aligned with pixel-level accuracy. This can potentially decrease segmentation model performance. In this paper, we highly customize the Blended Latent Diffusion Model (BLDM) [18] algorithm for generating the precisely located defect regions for industrial defect detection and localization.
Note that this work employs the diffusion model for generating synthetic datasets, just like the pioneering work [15, 43] do for more general purposes. Our method is orthogonal to those algorithms that use DPM to predict anomaly regions [44, 45, 46, 47].
III The proposed method
III-A Task Definition and Preliminaries
Following the conventional setting of anomaly detection, an anomaly-free (OK) sample set $\mathcal{X}_{OK}=\{\mathbf{x}^{i}_{OK}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}\mid i=1,2,\dots,N_{OK}\}$ is available for training. Furthermore, as introduced in Sec. I, a few defective (NG) samples $\mathcal{X}_{NG}=\{\mathbf{x}^{i}_{NG}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}\mid i=1,2,\dots,N_{NG}\}$ and the corresponding anomaly masks $\mathcal{M}_{NG}=\{\mathbf{M}^{i}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}\mid i=1,2,\dots,N_{NG}\}$ are given as the “seeds” of defect generation and usually $N_{NG}\ll N_{OK}$. The task of this paper is to generate the synthetic training set of defective samples as
[…]
where $\mathcal{X}^{\ast}_{NG}=\{\mathbf{x}^{\ast i}_{NG}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}\mid i=1,2,\dots,N^{\ast}_{NG}\}$ and $\mathcal{M}^{\ast}_{NG}=\{\mathbf{M}^{\ast i}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}\mid i=1,2,\dots,N^{\ast}_{NG}\}$ comprise the obtained synthetic sample images and the corresponding label masks, respectively. Furthermore, the defect generation algorithm usually sets $N^{\ast}_{NG}\gg N_{NG}$ to ensure higher sample variations.
In this paper, we propose to generate the synthetic samples using a diffusion model. It is common for diffusion models [15, 33] to approach generation target as a noise-free variable $\mathbf{x}_{0}$ and define the “forward diffusion process” as follows
$\begin{split}&q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\text{I})\\ &q(\mathbf{x}_{1:T}\mid\mathbf{x}_{0})=\prod^{T}_{t=1}q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}),\end{split}$ (2)
where $\{\beta_{t}\in(0,1)\}^{T}_{t=1}$ denotes the predefined variance schedule and I is a identity matrix. When $T$ is large enough, the diffusion process could convert a normal image into a random noise $\mathbf{x}_{T}$. On the contrary, the “reverse diffusion process” is defined as
$\begin{split}&p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{x}_{t},t))\\ &p_{\theta}(\mathbf{x}_{0:T})=p(\mathbf{x}_{T})\prod^{T}_{t=1}p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}),\end{split}$ (3)
where $\boldsymbol{\mu}_{\theta}(\cdot)$ and $\boldsymbol{\Sigma}_{\theta}(\cdot)$ are the two unknown functions that are usually estimated as deep models. In the test phase, given different random Gaussian noise samples $\mathbf{x}^{i}_{T},i=1,\dots,N_{tst}$, one could obtain high quality and various samples $\mathbf{x}^{i}_{0},\forall i$ (usually RGB images) from the same domain of the training samples. Akin to the above diffusion process, we could also generate lifelike anomalous images in $\mathcal{X}^{\ast}_{NG}$ based on a diffusion model which has been well learned on the samples sets $\mathcal{X}_{OK}$ and $\mathcal{X}_{NG}$.
III-B Latent Diffusion Model and Cross-Modal Prompts
Figure 2: The network structure of the proposed BLDM-based method for generating defective regions on a image. One can see besides the noise inupt, the model is governed by a a text prompt and a trimap that indicates the locations of the object and defect
III-B1 Latent diffusion model with controlling information
In this paper, we employ the Latent Diffusion Model (LDM) as the backbone of our method encouraged by its praiseworthy advantages [16]. As the name indicating, the forward diffusion and reverse diffusion processes of LDM are all performed in the latent space. In general, given the input image $\mathbf{x}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$, its corresponding latent feature $\mathbf{z}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$ is obtained using the encoder function $\Omega(\cdot)$ as $\mathbf{z}=\Omega(\mathbf{x})$ while the image can also be recovered via the decoder function $\Phi(\cdot)$ as $\mathbf{x}=\Phi(\mathbf{z})$. Based on the latent feature $\mathbf{z}$ and some simplifications  [^1] Readers are referred to [16] and [33] for details from Eq. 3, the reverse conditional probability in LDM is defined as
$\begin{split}&p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},t,\mathcal{C})=\\ &\mathcal{N}\left(\mathbf{z}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\right){\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathcal{C}),\sigma_{t}\text{I}\right),\end{split}$ (4)
where ${\epsilon}_{\theta}(\cdot)$ stands for the $\theta$-parameterized deep model to be trained; $\mathcal{C}$ denotes the controlling prompts fed into LDM for stabilizing the generation process [16]; $\alpha_{t}$, $\bar{\alpha}_{t}$ and $\sigma_{t}$ the parameters which can be calculated in a deterministic way at each $t$-th step. In our method, we propose to generate cross-modal controlling features to enhance the quality of the synthetic defect.
III-B2 Linguistic prompts
First, a linguistic prompt is obtained via changing the keywords of a following template sentence
[…]
where $obj$ and $def$ denote to the category of the current object (as the capsule in Fig. 2) and defect type (as the black contamination on the capsule), respectively. One can see this template design is straightforward and thus the prompt is always easy to obtain. In this work, the sentence $y$ is then processed by the parameter-frozen text encoder $\tau(\cdot):\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}^{d_{lang}}$ which is proposed in the CLIP algorithm [48]. In particular, as shown in Fig. 2, the linguistic controlling information is fed into the generation model $\epsilon_{\theta}(\cdot)$ and the generation process is influenced by the K-Q-V attention operations [16].
III-B3 Defect trimaps
Additionally, recall that in practical anomaly detection tasks, the defect-free samples are usually sufficient in terms of both amount and variation, we thus propose to generate defective samples under the guidance of the randomly selected defect-free samples. In a nutshell, a “trimap” $\Gamma\in\mathbb{R}^{H_{x}\times W_{x}}$ is designed to specify the desired location of the generated object and the defect, as shown in Fig. 2. To obtain a proper trimap, we firstly estimate the foreground region i.e., the interested object, by using the method proposed in [11] and store the foreground mask into $\mathbf{F}\in\mathbb{B}^{H_{x}\times W_{x}}$ with $\mathbf{F}(x,y)=1$ indicating a foreground at the coordinate $[x,y]$. Secondly, a synthetic defect mask is generated based on the genuine anomaly mask set $\mathcal{M}_{NG}$ as
[…]
where $\mathbf{M}^{i}_{NG},\mathbf{M}^{A}_{NG},\mathbf{M}^{\ast}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}$; “rand. $i$” refers to a random index selection; “rand. affine” denotes a random affine transform while “fit $\mathbf{F}$” stands for adjusting the location and scale of $\mathbf{M}^{A}_{NG}$ so that the adjusted mask is entirely inside the foreground region defined by $\mathbf{F}$. Finally, the value of each pixel on $\Gamma$ are defined as follows
$\Gamma(x,y)=\begin{cases}1\quad&\text{if}~{}\mathbf{M}^{\ast}_{NG}(x,y)=1,\\ 0.5\quad&\text{else if}~{}\mathbf{F}(x,y)=1,\\ 0\quad&\text{else}.\end{cases}$ (7)
As shown in Fig. 2, in the proposed AdaBLDM algorithm, the controlling trimap $\Gamma$ is firstly embedded by a convolutional block as $\zeta(\Gamma)\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$. The embedding feature is then fed into an encoder network $\hat{\epsilon}_{tri}(\cdot)$ that shares the same structure and initial parameters with the denoising encoder of $\epsilon_{\theta}(\cdot)$, except that the original “spatial attention” module is replaced with a self-attention process over the embedded trimap features. The deep features extracted from different layers of the denoising encoder are merged with the intermediate features of the denoising decoder of LDM via element-wise addition (see Fig. 2).
III-C Training Scheme
Formally, our cross-modal controlling features is defined as $\mathcal{C}=\{\tau(y),\Gamma\}$ and the learning objective of LDM is defined as:
$L_{LDM}\coloneqq\mathbb{E}_{\mathbf{z}_{t},\mathcal{C},\epsilon\thicksim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{C})\|^{2}_{2}\right].$ (8)
In this work, the denoising model ${\epsilon}_{\theta}(\cdot)$, the embedding block $\zeta(\cdot)$ of trimap prompts and the encoder $\hat{\epsilon}_{tri}$ for the latent trimap feature are all learned via optimizing the loss function defined in Eq. 8. However, different training schemes are applied for different modules. In specific, ${\epsilon}_{\theta}(\cdot)$ is firstly pre-trained by using the domain dataset (e.g. all the training images of all the subcategories of MVTec AD [19] for the experiments on this dataset) and then frozen during model learning for each specific subcategory; the latent feature encoder $\hat{\epsilon}_{tri}(\cdot)$ is initialized by the corresponding parameters of the original LDM model [16] and then fine-tuned for each subcategory based on the real defective samples; the convolutional block $\zeta(\cdot)$ is only learned based on the genuine anomaly samples with a random initialization.
Apart from the reverse diffusion model, the text prompts $y$ are obtained by using the CLIP [48] and BLIP [49] algorithm for sufficiently various descriptions. The text encoder is copied from CLIP and frozen during training. As to the image encoder $\Omega(\cdot)$ and decoder $\Phi(\cdot)$, we follow [16] to inherit the deep models of the VQ-VAE algorithm [50] and fix them during the whole training stage.
III-D Multi-Stage Denoising with Content Editing
Inspired the seminal work, namely Blended Latent Diffusion (BLD) [17], we propose to perform feature editing in the inference phase. However, due to the task difference, we extensively modify the editing scheme of BLD for defect generation. First of all, given a defect-free image $\mathbf{x}_{OK}$ and a trimap $\Gamma$ defined before, the corresponding defect mask $\mathbf{M}^{\ast}_{NG}$ is firstly calculated as introduced in Sec. III-B. Then the multi-stage denoising algorithm with content editing is performed and the workflow of this algorithm is illustrated in Alg. 1 where $\odot$ indicates element-wise multiplication and $\neg$ stands for the element-wise “Not” operation on the binary map.
Input: Defect-Free Image $\mathbf{x}_{OK}$, Trimap $\Gamma$, Defect Mask $\mathbf{M}^{\ast}_{NG}$, Language Prompt $y$, Decoder $\Psi(\cdot)$, Encoder $\Omega(\cdot)$, Range Numbers $T_{1}$, $T_{2}$ and $T_{3}$, Predefined Normal Distribution $\mathcal{N}(\mu_{z},\sigma^{2}_{z})$
Output: Blended Latent Feature $\mathbf{z}^{\ast}_{NG}$
1 $\mathbf{z}_{OK}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}\leftarrow\Omega(\mathbf{x}_{OK})$
2 $\mathbf{M}^{z}_{NG}\in\mathbb{B}^{H_{z}\times W_{z}}\leftarrow Dilate\&DownSample(\mathbf{M}^{\ast}_{NG})$
3 $\mathbf{z}_{t}\leftarrow RandSapmle(\mathcal{N}(\mu_{z},\sigma^{2}_{z}))$
4 $t=T_{1}+T_{2}+T_{3}$
5 /* Free Diffusion Stage */
6 while $t>T_{2}+T_{3}$ do
7       $\mathbf{z}_{t-1}\leftarrow Denoise(\mathbf{z}_{t},\Gamma,y)$
8       $t=t-1$
9 end while
10/* Latent Editing Stage */
11 while $t>T_{3}$ do
12       $\mathbf{z}_{t}=\mathbf{z}_{t}\odot\mathbf{M}^{z}_{NG}+\mathbf{z}_{OK}\odot\neg\mathbf{M}^{z}_{NG}$ $\mathbf{z}_{t-1}\leftarrow Denoise(\mathbf{z}_{t},\Gamma,y)$
13       $t=t-1$
14 end while
15/* Image Editing Stage */
16 while $t>=0$ do
17       $\mathbf{x}_{t}=\Psi(\mathbf{z}_{t})$
18       $\mathbf{z}_{t}\leftarrow\Omega(\mathbf{x}_{t}\odot\mathbf{M}_{NG}+\mathbf{x}_{OK}\odot\neg\mathbf{M}_{NG})$
19       $\mathbf{z}_{t-1}\leftarrow Denoise(\mathbf{z}_{t},\Gamma,y)$
20       $t=t-1$
21 end while
return $\mathbf{z}^{\ast}_{NG}=\mathbf{z}_{0}$
Algorithm 1 Multi-Stage Denoising with Editing
As shown in Fig. 2, the pipeline consistes of three stages in the denoising phase. The first stage is the conventional denosing process without content editing and takes $T_{1}$ steps. The second stage takes $T_{2}$ steps and the input feature $\mathbf{z}_{t}$ is merged with the feature of $\mathbf{x}_{OK}$, under the guidance of the dilated defect mask $\mathbf{M}^{\ast}_{NG}$. In the third stage, a novel content blending method proposed in this paper is conducted. In specific, the blending operation is performed in the pixel space rather than the latent space and the frozen encoder $\Omega(\cdot)$ and decoder $\Phi(\cdot)$ are used to switch the information between the two spaces.
Figure 3: The inference scheme of the proposed AdaBLDM algorithm. One can see that the whole procedure can be mainly divided into $4$ stages, namely the pure denoising stage without editing; the latent editing stage; the image editing stage, and the decoder adaptation stage.
III-E Online Decoder Adaptation
Recall that the aim of our algorithm is to generate better synthetic defective images which are actually reconstructed by the decoder $\Phi(\cdot)$, we thus propose to fine tune the decoder for each generated sample to achieve higher image quality.
Given the blended latent feature $\mathbf{z}^{\ast}_{NG}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$ obtained in Alg. 1, the demanding defect mask $\mathbf{M}^{\ast}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}$ and the source image $\mathbf{x}_{OK}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$, one can fine tune the decoder model $\Phi(\cdot)$ by conducting the online update algorithm described in Alg. 2
Input: Blended Latent Feature $\mathbf{z}^{\ast}_{NG}$, Defect-Free Image $\mathbf{x}_{OK}$, Maximum Step $T_{ft}$, Conservative Ratio $\lambda_{con}$, Anomaly Mask $\mathbf{M}^{\ast}_{NG}$, Decoder $\Phi(\cdot)$
Output: Refined Image $\mathbf{x}^{\ast}_{NG}$
1 $\mathbf{x}^{\ast}_{NG}\leftarrow\Phi(\mathbf{z}^{\ast}_{NG}),~{}\bar{\mathbf{M}}^{\ast}_{NG}(i,j)\leftarrow\neg\mathbf{M}^{\ast}_{NG}(i,j)$
2 $t=1$
3 while $t<=T_{ft}$ do
4       $\tilde{\mathbf{x}}\leftarrow\Phi(\mathbf{z}^{\ast}_{NG})$
5       $L_{i}=\sum_{i,j}^{H_{x}W_{x}}\left(\tilde{\mathbf{x}}(i,j)-\mathbf{x}_{OK}(i,j)\right)^{2}\odot\bar{\mathbf{M}}^{\ast}_{NG}(i,j)$
6       $L_{d}=\sum_{i,j}^{H_{x}W_{x}}\left(\tilde{\mathbf{x}}(i,j)-\mathbf{x}^{\ast}_{NG}(i,j)\right)^{2}\odot\mathbf{M}^{\ast}_{NG}(i,j)$
7       $\Phi(\cdot)\leftarrow AdamW(L_{i}+\lambda_{con}L_{d},\Phi(\cdot))$
8       $t=t+1$
9 end while
[…]
return $\mathbf{x}^{\ast}_{NG}$
Algorithm 2 Online Decoder Adaptation
One can see that the online adaptation strategy encourages the decoder $\Phi(\cdot)$ generating similar defect-free pixels as $\mathbf{x}_{OK}$ while outputting the defective pixels as the original decoder suggests. The conservative ratio $\lambda_{con}$ balance the two requirements. In the experiments of this work, we prove that the individually updated decoder can yield better samples and lead to higher AD performance.
III-F Implementation Details
In this paper, all the images are resize to $256\times 256$ following the setting of DeSTseg [6] which is a SOTA anomaly detection algorithm and adopted by this paper for evaluating the quality of generated samples. In this way, we set $H_{x}=W_{x}=256$ and $H_{z}=W_{z}=32$. The latent feature channel number $C_{z}$ is $4$ in this paper for efficient denoising operation. As to the trimap prompts, the object’s region is estimated by using the foreground estimation method proposed in [11] and the anomalous mask is obtained based on the randomly selected $10$ genuine defective samples for each subcategory of the dataset. The affine transformation matrix mentioned in Sec. III-B3 is determined by a random rotation angle $\gamma\in[0^{\circ},360^{\circ}]$ and a random scaling factor $s\in[0.85,1.15]$. For the AD tasks on texture subcategories, such as a wooden or a fabric surface, all the pixels belong to the foreground region except the defective ones.
We employ the AdamW optimizer [51] for updating the model parameters, with a default learning rate of $1\times 10^{-5}$, and default betas are set to $0.9$ and $0.999$, respectively. During training, there is a $10\%$ chance that empty text prompt is input to the model.
During the inference stage, we adopt the DDIM [34] sampling method with the normal denoising step $T_{1}=50$, the latent blended step $T_{2}=30$, and pixel blended step $T_{3}=5$. In the online decoder adaptation stage, $T_{ft}=200$ and the conservative ratio $\lambda_{con}=100$ to avoid overfitting to the anomaly-free region. The same optimizer as offline learning is used for this online adaptation except the learning rate is set to $1\times 10^{-4}$.
IV Experiments
IV-A Experiment Setting
IV-A1 Evaluation Methodology
In this section, we evaluate the proposed method for generating defective samples practically and extensively. In most AIGC applications, the appearance quality of the generated image, measured in either quantitative or qualitative ways, is the most important criterion to evaluate a generation algorithm. However, in the scenario of defective sample generation, we claim that a more important measuring standard is the improvement in AD accuracy. Consequently, a state-of-the-art anomaly detection algorithm, namely DeSTseg [6] is learned based on the augmented dataset, and the corresponding performance gain (if any) directly reflects the effectiveness of the synthetic samples. In addition, the classical Support Vector Machine (SVM) [52] is also employed as a complement to DeSTseg. Specifically, the SVM model is learned to classify the pixels into normal and anomalous, and in this way, a vanilla anomaly detection/localization method is obtained. The combination of the sophisticated (DeSTseg) and the naive (SVM) AD algorithms increases the comprehensiveness and thus the reliability of our experiment.
IV-A2 Comparing Generation Methods
We compare our defect generation algorithm with the cutting-edge generative models including DFM [14] and DCDGANc [13] which is built based on StarGAN [53] or StyleGANv2 [29] algorithms. However, as the original StyelGANv2 in DCDGANc can easily lead to training crash according to our experiment, in this work, the StyleGANv2 is replaced by StyleGANv2-Ada [54] to ensure the stability of training. Furthermore, the results of the original DeSTseg algorithm, which also generates synthetic defects in the pasting fashion, are also compared. The training samples of the compared methods are strictly identical to those of AdaBLDM.
IV-A3 Datasets
To achieve a more general study, we perform the comparison on three well-known datasets, namely MVTEC-AD [19] ($7$ subcategories over $15$), BTAD [55] (full dataset), and KSDD2 [56] which are introduced below.
  • MVTec AD[19] is an open dataset containing ten object categories and five texture categories commonly seen, with up to eight defect categories for each object/texture category. All the images are accompanied by pixel-level masks showing the defect regions. Among the available defect categories, we opted to utilize defect categories that contain more than $10$ defect images as the training dataset for our model. Specifically, we randomly selected $10$ defect images per defect category from the test set and total defect-free images from the train set for model training and defect data generation. Due to time constraints, within the MVTec AD dataset, we focused on evaluating our model using data from $2$ Object categories (Hazelnut and Capsule) and $5$ Texture categories.
  • BTAD [55] comprises $2540$ images from three categories of real-world industrial products with different body and surface defects. It is usually considered as a complementary dataset to the MVTec AD when evaluating an AD algorithm. Consistent with MVTec AD, we similarly opt for the selection of 10 instances of defect data along with defect-free data for the generation and training of the model.
  • KSDD2 [56] with over $3000$ images obtained while addressing a real-world industrial problem. This dataset comprises only one category, focusing on surface defects. For model training and data generation, we exclusively utilize the defect data and defect-free data from the training set.
To the best of our knowledge, the experiment of this paper involves much more subcategories, both in terms of number and variations, than any others in the current literature.
IV-A4 Evaluation metrics for Anomaly Detection and Localization
  • Pixel-AUC is the area under the receiver operating characteristic curve at the pixel level. It is the most popular AD measuring method while failing to reflect the real performance difference between algorithms when a serious class imbalance exists.
  • PRO [57] on the contrary, focuses on the anomaly pixels and treats the AD performance on each anomaly region equally. Consequently, the PRO metric is more robust to the class imbalance which is a common situation in most AD benchmarks.
  • AP [58] metric as a conventional metric for semantic segmentation, is frequently adopted in recently proposed AD algorithms. It reflects the anomaly detection performance from a pixel-level perspective.
  • IAP [6] focus on instance recall as a more straightforward metric.
  • IAP90 For those applications requiring an extremely high recall, the precision at recall = kis also computed and denoted as IAP@k, and k we follow [6] by setting k = $90$ in our experiments.
IV-A5 Evaluation metrics for generation quality
Following the conventional evaluation manner for image generation algorithms, we here also compare the scores related to the quality of the generated images. In particular, the Kernel Inception Distance (KID) [59] as well as the Learned Perceptual Image Patch Similarity (LPIPS) [60] are involved in the comparison. In simple words, the former criterion reflects the generation’s authenticity while the latter one indicates the generation’s diversity.
IV-A6 Training Settings for the Involved AD Models
In this paper, we propose to evaluate the generated samples by using the simplest AD method and a sophisticated AD method. In this way, one can obtain a thorough examination of the generated samples.
As one end of the spectrum, the SVM-based anomaly detection algorithm is extremely simple and thus can directly reflect the patch-wise quality of the generated images. In specific, the SVM model based on the RBF kernel [61] is employed as the patch classifier. The patch features are extracted by using the PatchCore algorithm [4] and each $768$-D feature is assigned a positive label if the corresponding patch is normal while negative labels are assigned to the anomalous ones. In each trial, $100$ images are randomly sampled from the synthetic defective images and then $5000$ positive patch features and $5000$ negative patch features are extracted for training the SVM model. The final performance on each subcategory is calculated by averaging the results of $10$ random trails with the same setting. Besides, the performance of the SVM model trained merely on the genuine defective data is also reported. As to the parameter setting, we by default set the $\gamma$ of the RBF kernel as $1\times 10^{-4}$ while the regularization parameter $C$ is set to $1$.
As a SOTA anomaly detection method, DeSTseg [6] employs a discriminative model to detect anomalous regions based on comparing them and their reconstructed defect-free counterparts. Before training, we first generate $5000$ defective samples by using AdaBLDM and another $5000$ defective samples by using the data augmentation method of DeSTseg. We store the defective samples generated by AdaBLDM in $\mathcal{D}_{gen}$ and store those augmented by DeSTseg in $\mathcal{D}_{raw}$. The training data is then sampled from the defect-free samples, $\mathcal{D}_{raw}$ and $\mathcal{D}_{gen}$ in the ratio of $5:4:1$. Note that according to the methodology of DeSTseg, the genuine defective samples, which are the seeds of the synthetic data generation, can not be used for training as the anomaly-free version of them is unknown. For each subcategory, the DeSTseg is learned for $10000$ iterations and we pick the DeSTseg model with a UNIQUE iteration number for all the subcategories in one dataset.
IV-B Quantitative Results
IV-B1 Results on MVTec AD
We first conduct the comparison of the SVM-based anomaly detection on the well-known MVTec AD dataset and report the results in Tab. I. From the table, we can see that the SVM achieves consistently best accuracies based on the samples generated by AdaBLDM. The genuine samples achieve the second-best performance while the samples obtained by comparing SOTA generation algorithms perform worse than the two aforementioned competitors.
TABLE I: The Anomaly Detection and Localization Performance of SVM on the MVTec AD dataset, Pixel-Auc/PRO/AP/IAP/IAP90, the best and the second-best numbers are shown in red and blue, respectively.
Category Genuine DFM[14] (AAAI2023) DCDGANc-StarGAN[13] (CVPRW2023) DCDGANc-StyleGAN[13] (CVPRW2023) AdaBLDM (ours)
Hazelnut 98.07/70.03/47.07/94.85/63.03 97.65/66.05/38.7/93.87/57.6 82.41/54.24/19.19/71.69/52.3 96.79/67.63/41.27/94.18/59.98 98.23/71.08/47.94/94.88/66.65
Wood 95.22/63.53/36.95/91.22/55.33 39.91/3.9/2.89/12.99/5.12 96.79/67.63/41.27/94.18/59.98 89.83/51.97/25.08/81.6/44.96 95.96/71.02/48.83/93.05/63.41
Capsule 92.28/13.91/6.65/84.01/12.31 91.72/37.37/9.01/87.36/25.69 99.51/71.09/52.11/99.14/61.98 94.05/37.13/10.47/89.09/28.15 94.17/45.26/13.12/91.77/32.86
Leather 99.07/62.88/40.86/98.52/51.21 45.15/3.05/0.66/29.62/2.25 89.83/51.97/25.08/81.6/44.96 99.51/71.09/52.11/99.14/61.98 99.23/68.5/60.52/99.56/61.49
Grid 89.02/28.68/2.87/75.29/23.78 73.75/3.53/0.78/47.6/3.7 93.66/46.29/19.09/83.09/45.71 95.82/46.91/12.81/92.49/42.47 95.22/35.31/7.55/89.01/32.86
Tile 93.94/74.81/49.55/87.72/68.71 73.54/38.69/7.85/50.4/43.48 94.05/37.13/10.47/89.09/28.15 82.41/54.24/19.19/71.69/52.3 92.74/71.86/47.4/85.9/68.46
Carpet 95.22/63.53/36.95/91.22/55.33 39.91/3.9/2.89/12.99/5.12 95.54/44.8/13.12/91.49/40.75 89.83/51.97/25.08/81.6/44.96 95.96/71.02/48.83/93.05/63.41
Average 95.06/54.48/32.97/89.49/48.11 72.09/24.93/9.76/54.7/22.52 93.11/53.31/25.76/87.18/47.69 93.23/54.0/26.34/87.62/48.1 95.95/57.69/35.51/91.57/52.53
On the other hand, Tab. II illustrates the AD performances of DeSTseg based on the MVTec AD dataset. As shown in the table, the proposed method ranks the first for all the $5$ AD metrics. In particular, the performance gains brought by the synthetic samples generated by AdaBLDM are $1.00\%$, $1.98\%$ $6.68\%$, $6.35\%$ and $7.93\%$ on Pixel-AUC, PRO, AP, IAP, and IAP90, respectively. The significant improvements prove the validness of the proposed method.
TABLE II: The Anomaly Detection and Localization Performance of DeSTseg on the MVTec AD dataset, Pixel-Auc/PRO/AP/IAP/IAP90, the best and the second-best numbers are shown in red and blue, respectively.
Category DESTSEG[6] (CVPR2023) DFM[14] (AAAI2023) DCDGANc-StarGAN[13] (CVPRW2023) DCDGANc-StyleGAN[13] (CVPRW2023) AdaBLDM (ours)
Hazelnut 99.65/97.46/88.22/89.06/75.58 99.45/98.01/88.5/90.48/82.32 99.53/97.69/87.47/87.78/74.17 99.62/98.1/90.31/90.69/80.78 99.65/98.65/91.27/91.47/81.96
Wood 95.72/93.0/73.62/79.77/63.6 97.91/97.09/83.78/92.49/85.63 97.83/96.99/84.14/92.63/85.05 98.34/97.17/84.67/91.98/85.24 98.23/97.23/84.59/90.85/82.47
Capsule 99.21/96.29/57.79/52.7/33.32 98.43/95.39/62.89/58.5/31.82 98.44/94.78/61.07/56.63/30.49 98.85/95.77/62.08/57.28/34.16 98.61/96.14/63.39/61.04/39.0
Leather 99.76/99.06/75.0/74.96/64.92 99.81/99.42/80.08/83.28/72.91 99.84/99.56/82.53/86.47/76.82 99.84/99.5/81.7/83.64/74.32 99.84/99.56/82.62/85.89/73.51
Grid 98.82/94.3/57.74/57.11/35.25 98.24/94.38/56.0/57.38/26.34 98.21/93.52/57.2/57.52/20.91 97.88/91.47/56.4/55.95/13.72 98.95/96.21/60.65/62.87/31.11
Tile 98.7/97.28/93.12/96.38/92.33 98.68/96.77/92.87/95.58/88.65 99.04/97.54/94.1/96.41/90.38 99.1/97.69/93.97/96.27/90.24 99.17/97.86/94.57/96.61/90.14
Carpet 94.86/92.99/69.74/84.79/58.4 97.89/96.64/78.99/88.6/74.26 98.18/97.08/81.45/89.28/76.53 98.55/97.5/82.07/90.34/78.81 99.33/98.47/84.87/90.53/80.77
Average 98.1/95.77/73.6/76.4/60.49 98.63/96.81/77.59/80.9/65.99 98.72/96.73/78.28/80.96/64.91 98.88/96.74/78.74/80.88/65.32 99.11/97.73/80.28/82.75/68.42
IV-B2 Results on BTAD and KSDD2
Similarly, we compare the proposed method with other generation algorithms in terms of the AD performance on the other two well-known AD datasets. The accuracies obtained by DeSTseg are summarized in Tab. III and Tab. IV for BTAD [55] and KSDD2 [56], respectively. One can again observe the remarkable superiority of the proposed method in the two tables. The artificial defects simulated by AdaBLDM lead to the best performances for all the involved AD metrics, on both two datasets.
TABLE III: The Anomaly Detection and Localization Performance of DeSTseg on the BTAD dataset, Pixel-Auc/PRO/AP/IAP/IAP90, the best and the second-best numbers are shown in red and blue, respectively.
Category DESTSEG[6] (CVPR2023) DFM[14] (AAAI2023) DCDGANc-StarGAN[13] (CVPRW2023) DCDGANc-StyleGAN[13] (CVPRW2023) AdaBLDM (ours)
01 96.25/82.5/45.74/39.93/26.95 95.97/74.17/37.13/31.71/14.92 95.33/71.39/35.37/31.2/11.32 95.64/82.34/38.15/36.24/26.78 96.84/84.25/43.22/39.04/27.93
02 95.73/57.99/60.58/40.82/7.96 96.15/64.78/62.73/47.46/12.61 94.67/60.65/55.6/49.6/10.63 94.39/61.32/64.72/45.09/8.46 96.55/65.46/70.16/47.99/13.24
03 99.47/98.01/44.29/35.87/17.54 98.94/95.45/39.13/34.18/19.22 98.92/95.32/37.26/35.82/18.02 99.27/97.43/43.5/36.58/16.2 99.47/97.97/45.91/36.59/17.95
Average 97.15/79.5/50.2/38.87/17.48 97.02/78.13/46.33/37.78/15.58 96.31/75.79/42.74/38.87/13.32 96.43/80.36/48.79/39.3/17.15 97.62/82.56/53.1/41.21/19.71
TABLE IV: The Anomaly Detection and Localization Performance of DeSTseg on the KSDD2 dataset, Pixel-Auc/PRO/AP/IAP/IAP90, the best and the second-best numbers are shown in red and blue, respectively.
Method Pixel AUC PRO AP IAP IAP90
DESTSEG[6] (CVPR2023) 89.45 83.42 52.3 57.59 1.29
DFM[14] (AAAI2023) 86.69 85.53 47.38 54.39 1.26
DCDGANc-StarGAN[13] (CVPRW2023) 93.08 85.3 46.53 49.96 1.69
DCDGANc-StyleGAN[13] (CVPRW2023) 92.38 84.83 51.72 55.65 1.56
AdaBLDM (ours) 96.42 92.93 70.62 73.88 7.6
IV-B3 Image Quality
Tab. V demonstrates the scores on the image quality of the comparing generation algorithms. From the table, we can see that the proposed method achieves the best KID performances for most subcategories (except Grid and Wood) as well as the overall comparison. This is consistent with the AD performances reported in Tab. II as the KID metric reflects the “reality” of the generated images.
TABLE V: The KID×$\times 10^{3}$@5k $\downarrow$ and LPIPS@5k $\uparrow$ of the generated images on MVTec AD, the best and the second-best numbers are shown in red and blue, respectively
Category DFM[14] (AAAI2023) DCDGANc-StarGAN[13] (CVPRW2023) DCDGANc-StyleGAN[13] (CVPRW2023) AdaBLDM (ours)
Hazelnut 4.74/0.188 1.76/0.16 1.97/0.167 1.72/0.182
Wood 48.47/0.172 7.91/0.046 8.74/0.048 2.74/0.076
Capsule 3.66/0.018 1.87/0.026 1.08/0.026 4.14/0.027
Leather 61.35/0.199 9.23/0.088 13.6/0.078 5.85/0.032
Grid 4.00/0.061 2.04/0.073 7.21/0.087 5.38/0.07
Tile 16.74/0.015 23.16/0.036 21.5/0.05 12.83/0.046
Carpet 3.64/0.082 4.29/0.095 3.61/0.088 3.55/0.073
Average 20.37/0.105 7.18/0.075 8.25/0.078 5.17/0.072
In contrast, it is interesting to see that our method lead to poor LPIPS scores compared with other generation algorithms. Recall that either the editing process in the denoising stage or the online decoder adaptation prevents the generated image from being too diverse from the source image and the desired defect. In this way, we claim that the poor image diversity might be a disadvantage of the proposed algorithm from the traditional perspective while it also implies a better-controlled generation process that can benefit the following training stage of AD models.
IV-C Qualitative Results
To offer readers a more direct impression of the proposed image generation algorithm, We hereby depict the generated defects by using the comparing methods in Fig. 4. One can see the proposed method, either with or without the online adaptation stage, can achieve significantly more lifelike images compared with other methods. In particular, only our method can generate the “crack” defects on the hazelnut subcategory. On the contrary, the DFM algorithm, though can generate synthetic defective samples with good quality in some scenarios, could fail in generation for some other cases.
Figure 4: The genuine and synthetic defective samples on MVTec AD dataset. From top to bottom, the samples with genuine defects; the samples generated by using DCDGANc-StarGAN; the samples generated by using DCDGANc-StyleGAN; the samples generated with DFM and the samples generated by using our method.
We further illustrate the comparison in a more fine-grained level in Fig. 5. From the figure we can see that the proposed method is more well-controlled in generating the defect regions with the specific category and location.
Figure 5: The fine-grained comparison in the generation quality on the Hazelnut subcategory of MVTec AD. Different kinds of defects are mimicked by the involved generation algorithms and their performances can be compared on a more detailed level.
In addition, the generated images with synthetic defects on BTAD and KSDD2 are also shown in Fig. 6. One can see that besides MVTec AD, the image quality of AdaBLDM is maintained for the two different datasets.
Figure 6: The generation outputs of AdaBLDM on BTAD and KSDDV2 datasests.
Finally, as a qualitative complement to Tab. II, Fig. 7 shows the heat map of anomaly predicted by the DeSTseg algorithm. We can see that thanks to the higher quality of the simulated synthetics, the AdaBLDM-equipped AD algorithm can output the best heat maps for anomaly detection and localization.
Figure 7: Test results of different DeSTseg models learned based on the training sets obtained by using different generation methods. All the involved subcategories in this work ($7$ from MVTec AD, $3$ from BTAD, and the only one in KSDDV2) are shown here.
IV-D Ablation Study
To understand the success of the proposed AdaBLDM algorithm, we utilize ablation experiments to evaluate the specific contributions of the primary components which include:
  • Pretrain with domain knowledge
    As introduced in Sec. III-C, the denoising model ${\epsilon}_{\theta}(\cdot)$ is pretrained on the domain data. If this module is removed, AdaBLDM will rely on the prior knowledge from the LAION-5B dataset [62], which might not highly correlate with the domain of industrial AD. Note that the pretraining is strictly independent to the current subcategory to maintain the fair comparison.
  • Image Editing
    Different from the conventional blended latent diffusion model, we propose to perform the intent editing in both latent space and pixel space (see Alg. 1). If this module is removed, the editing is only performed in the latent space.
  • Online Decoder Adaptation
    The online decoder adaptation is introduced in Alg. 2. The absence of this module results in a fixed image decoder.
The ablation results are summarized in Tab. VI. It is clear that as each module mentioned above is added to the algorithm, the corresponding AD performance increases steadily. In particular, the total performance improvements are $9.78\%$ in AP, $14.46\%$ in IAP, and $13.98\%$ in IAP90, respectively.
TABLE VI: Ablation Study of AdaBLDM
Modules Performance
Pretrain Image Editing Adaptation Pixel auc Pro AP IAP IAP90
96.51 92.18 70.50 68.29 54.44
97.23 93.14 71.21 69.30 58.32
99.12 97.49 80.25 82.36 67.65
99.11 97.73 80.28 82.75 68.42
V Conclusion & future work
In this paper, we propose a sophisticated method to generate realistic defective samples for industrial anomaly detection and localization. Motivated by the great success of the diffusion model, we highly customize the Blended Latent Diffusion Model (BLDM) with several novel modifications for this specific task. The yielded algorithm is termed AdaBLDM and can generate better synthetic samples compared with the existing methods based on conventional “cut-and-paste” scheme or Generative Adversarial Networks (GANs). By applying the AdaBLDM for AD data augmentation, one can observe significant increase on the AD performances of SOTA algorithms. The AIGC algorithm proposed in this paper can be directly used as an effective data augmentation method in real-world applications and also sheds light on the influential factors for solving defect generalization task. In the future, one could obtain even better synthetic defective samples by introducing finer controlling signals such as the language prompts describing the details (shape, size etc.) about the demanding defects. Another appealing direction of future work is to speed up the denoising process of the diffusion model so that more defective samples can be generated given a limited time budget, which is a crucial constraint in most practical scenarios.

### v2

@@
 The proposed method tailors the blended latent diffusion model for defect sample generation, employing a diffusion model to generate defective samples in the latent space.
-A feature editing process, controlled by a ”trimap” mask and text prompts, refines the generated samples.
+A feature editing process, controlled by a “trimap” mask and text prompts, refines the generated samples.
 The image generation inference process is structured into three stages: a free diffusion stage, an editing diffusion stage, and an online decoder adaptation stage.
@@
 In a practical manufacturing workflow, obtaining defective samples is considerably more challenging compared to acquiring defect-free ones.
-Consequently, the majority of recently proposed algorithms for industrial defect inspection, such as [1, 2, 3, 4, 5, 6, 7], address the issue as an Anomaly Detection (AD) problem—an established machine learning challenge [8, 9].
-In these industrial AD algorithms, the assumption is made that all anomaly-free samples, be they complete images or image patches, belong to a single distribution.
+Consequently, the majority of recently proposed algorithms for industrial defect inspection, such as [1, 2, 3, 4, 5, 6, 7, 8, 9, 10], address the issue as an Anomaly Detection (AD) problem—an established machine learning challenge [11, 12].
+Among these industrial AD algorithms, the assumption is made that all anomaly-free samples, be they complete images or image patches, belong to a single distribution.
 Meanwhile, defective samples are identified as “outliers”.
@@
 Figure 1: Illustration of three defect generation styles. From top to bottom: conventional approaches, GAN-based algorithms, and the proposed method.
-On the flip side, achieving this compatibility comes at the expense of an extremely imbalanced data distribution, a characteristic that is deemed unacceptable in the context of most discriminative algorithms.
+Nonetheless, achieving this compatibility comes at the expense of an extremely imbalanced data distribution, a characteristic that is deemed unacceptable in the context of most discriminative algorithms.
 Consequently, some researchers suggest the generation of artificial defect patterns using conventional approaches [1, 3, 5] to facilitate subsequent discriminative learning processes.
-More recently, state-of-the-art (SOTA) algorithms [10, 11, 12] highlight that, in addition to synthetic defects, incorporating a few real anomalous samples during the training stage can significantly enhance Anomaly Detection (AD) accuracy.
-This approach holds the promise of achieving even higher performance when superior simulated defective patterns are integrated into training.
-Pioneering works [13, 14] have already demonstrated the benefits of generating these ”lifelike” defects using more sophisticated methods.
+Recently, state-of-the-art (SOTA) algorithms [13, 14, 15] suggest that incorporating a few real anomalous samples during the training stage can significantly enhance the AD accuracy if these samples are combined with synthetic samples.
+For example, some pioneering works [16, 17] demonstrated the benefits of generating these ”lifelike” defects using more sophisticated methods.
 The first two rows of Fig.
-1 illustrate the existing two styles of defect generation.
-In this paper, we propose the integration of the cutting-edge AI generation algorithm, namely, the Diffusion Probabilistic Model (DPM) [15], into the realm of defect generation.
-Specifically, we leverage the Blended Latent Diffusion Model [16, 17, 18] as our foundational framework.
-To enhance the generation of defective samples, we implement three key modifications to the original Blended Latent Diffusion Model (BLDM).
-Firstly, we design a novel “defect trimap” to delineate the target object mask and the defective regions on the generated image, incorporating it as a new form of controlling information alongside the language prompts within the stable diffusion model.
-Secondly, to ensure the authenticity of the generated samples, a cascade “editing” stage is introduced for the BLDM, operating in both the latent and pixel space.
+1 illustrate two different types of defect generation.
+In this paper, we propose the integration of the cutting-edge AI generation algorithm, namely the Diffusion Probabilistic Model (DPM) [18], into the realm of defect generation.
+Specifically, we leverage the Blended Latent Diffusion Model [19, 20, 21] to create our framework.
+To enhance the generation of defective samples, we innovate the Blended Latent Diffusion Model (BLDM)[21] with three modules.
+First, we design a novel “defect trimap” to delineate the target object mask and the defective regions on the generated image, incorporating it as a new form of controlling information alongside the language prompts within the stable diffusion model.
+Second, to ensure the authenticity of the generated samples, a cascaded “editing” stage is introduced into the BLDM, at both the latent and pixel space.
 Finally, an innovative online adaptation of the image encoder is proposed to refine the quality of the generated images.
-Consequently, the highly customized BLDM algorithm, referred to as AdaBLDM in this paper, exhibits the ability to generate defective samples with significantly higher image quality and pattern variations compared to those produced by state-of-the-art (SOTA) methods.
+Consequently, the customized BLDM algorithm, referred to as AdaBLDM in this paper, is capable of generating higher-quality defective samples with pattern variations.
 The third row of Fig.
@@
 II-A Synthetic Defect Generation
-As introduced in Sec.
-I, due to a scarcity of defective samples, many researchers [1, 2, 3, 5, 6] are “compelled” to generate synthetic defective samples for better AD performance.
-The most straightforward way to mimic defects is to “paste” anomaly pixels on the normal images as a special case of data augmentation.
-Baseline methods like [20], CutPaste [21], and [1] randomly cut regions from normal images and paste them to the “incorrect” places as artificial defects.
+Due to the scarcity of defective samples, generating synthetic defective samples becomes a rule of thumb among researchers to achieve satisfactory performance [1, 2, 3, 5, 6, 23].
+A straightforward way akin to data augmentation is mimicking the defects by “pasting” anomalous pixels on the normal images.
+For example, methods in the regime [24, 25, 1] randomly cut regions from normal images and paste them to the “incorrect” places as artificial defects.
 Crop&Paste [26] and PRN[27] crop defect areas from genuine defective images and paste them onto defect-free images.
@@
 However, in these cases, the distribution of fabricated defects differs from the real one and thus the performance gain can not be ensured.
-Encouraged by the great success of AIGC (AI Generated Contents) methods including GAN (Generative Adversarial Network) algorithms [25] and Diffusion Models [26], some efforts have been made to generate more natural-looking and various simulated defects.
+Inspired by the potent AIGC (AI Generated Contents) methods such as Generative Adversarial Network (GAN) based algorithms [29, 30, 31] and diffusion model-based algorithms [32], some attempts have been devoted to generating more natural-looking simulated defects.
 In particular, SDGAN [30] employs two generators to switch the defective and non-defective status for sample images.
-SDGAN can generate a high-quality, diverse dataset of steel surface with both defective and defect-free images, thus higher classification performance is obtained.
-Similarly, Defect-GAN [28] can convert non-defective images into defective ones to enhance the learning process of the defect classifier.
-However, the above two GAN-based generation algorithms can not generate well-defined defective regions.
-Consequently, they could boost the performance of defect classification while can hardly lift the defect segmentation accuracy which is crucial in many real-world applications.
-To the best of our knowledge, [13] and [14] are the only two exiting works that generate both the synthetic defective images as well as the corresponding pixel-level labels.
-In specific, the DCDGANc algorithm proposed in [13] design a novel GAN-based method to mimic real defects and the defects are fused with defect-free images via an improved Poisson blending algorithm.
-On the other hand, [14] introduced DFMGAN, stems from StyleGANv2 [29], to generate defect images and defect masks by using the proposed defect-aware residual blocks.
-Although the above two GAN-based methods can generate high-quality images with defects, they fall short in generating the defect mask strictly aligned to the generated defective pixels on the image.
-The misalignment stems from the disparate nature of generative methods which is a blessing in image diversity whereas it could significantly mislead the defect segmentation model in the scenario of AD.
+As a result, SDGAN [30] can generate a high-quality, diverse dataset of steel surfaces with both defective and defect-free images.
+In a similar vein to SDGAN [30], Defect-GAN [33] converts non-defective images into defective ones to enhance the learning process of the defect classifier.
+However, the above two GAN-based generation algorithms cannot generate defective regions that are aligned precisely as required.
+Despite that they could enhance the performance of defect classification, they cannot lift the defect segmentation accuracy which is crucial in many real-world applications.
+More recently, [16] and [17] proposed to generate both synthetic defective images and the corresponding pixel-level labels.
+In specific, the DCDGANc algorithm [16] designed a novel GAN-based method to mimic real defects, which can be fused with defect-free images via an improved Poisson blending algorithm.
+At the same time, [17] introduced DFMGAN, stemming from StyleGANv2 [34], to generate defective images and defective masks by using the proposed defect-aware residual blocks.
+Although the above two GAN-based methods can generate high-quality images with defects, they fall short in generating the defective mass strictly aligned to the generated defective pixels on the image.
+Such misalignment is ascribed to the disparity between the image-specific generative methods and the challenging scenario of AD, and thus significantly misleads the defective segmentation model in AD.
 In contrast, in this paper, we propose a novel defect generation method based on Diffusion Models.
-We tame this AIGC algorithm to generate diverse, realistic but accurately-controlled defective image samples via a series of novel modifications.
+We tame this general-purpose algorithm to generate diverse, realistic but accurately-controlled defective image samples via a series of novel modifications.
 II-B Diffusion Probabilistic Models for Image Editing
-Compared with the more traditional generation methods like Generative Adversarial Networks (GANs) [30], Variational AutoEncoders (VAEs) [31] and flow-based algorithms [32], Diffusion Probabilistic Models (DPMs) [15, 33, 34, 35] enjoy a higher quality of the synthetic images and impressive stability of training.
-Accordingly, during the past few years, DPM-based approaches, such as Stable Diffusion (Stability AI) [36], DALL-E2 (OpenAI) [37], and Imagen (Google) [38], have achieved state-of-the-art performances in the task of content generation.
-In the subdomain of image editing, DPM-based methods also play the most important role nowadays.
-In particular, the Latent Diffusion Model [16] conducts the image generation or editing in the latent space with a much less feature dimensionality and thus achieve much higher editing efficiency.
-RePaint [39] fully leverages a pre-trained DDPM [34] to perform image editing by sampling given pixels during the reverse diffusion stage.
-DCFace [40] employs two feature encoders to control the Diffusion Model, facilitating facial image editing.
-[41] proposes the ControlNet to govern existing Diffusion models to achieve a better-controlled synthesis.
-[17, 18] enables image editing by making modifications in either the pixel space or feature space.
-[42] introduces sparse text optimization to achieve better image editing.
-Those methods, though achieve SOTA performance in general image editing tasks, the fuzzily governed synthesis easily fails to generate the image regions aligned with pixel-level accuracy.
-This can potentially decrease segmentation model performance.
-In this paper, we highly customize the Blended Latent Diffusion Model (BLDM) [18] algorithm for generating the precisely located defect regions for industrial defect detection and localization.
-Note that this work employs the diffusion model for generating synthetic datasets, just like the pioneering work [15, 43] do for more general purposes.
-Our method is orthogonal to those algorithms that use DPM to predict anomaly regions [44, 45, 46, 47].
+In the last few years, researchers have proposed various approaches to realize high-quality image generation.
+The off-the-shelf tools include Generative Adversarial Networks (GANs) [35], Variational AutoEncoders (VAEs) [36], the flow-based algorithms [37], and Diffusion Probabilistic Models (DPMs) [18, 38, 39, 40].
+In specific, DPMs have demonstrated higher qualities of the synthetic images and impressive stability of training.
+Accordingly, the DPM-based approaches, such as Stable Diffusion (Stability AI) [19], DALL-E2 (OpenAI) [41], and Imagen (Google) [42], have achieved state-of-the-art performances in the task of content generation.
+DPM-based methods also play an important role in the remit of image editing.
+In particular, the Latent Diffusion Model [19] conducts the image generation or editing in the lower-dimensional latent space and thus can achieve higher efficiency; RePaint [43] fully leverages a pre-trained DDPM [39] to perform image editing by sampling given pixels during the reverse diffusion stage; DCFace [44] employs two feature encoders to control the Diffusion Model to facilitate facial image editing; [45] proposes the “ControlNet” to govern existing Diffusion models to achieve a better-controlled synthesis; [20, 21] enable image editing by making modifications in either the pixel space or feature space; [46] propose the ”Dual-Cycle Diffusion” to generate an unbiased mask to guide image editing.
+While these methods demonstrate state-of-the-art (SOTA) performances in general image editing tasks, they can hardly produce image regions that are precisely aligned at the pixel level due to the inherently fuzzy nature of the reverse diffusion process.
+This nuisance can potentially decrease the segmentation capacity in anomaly detection.
+In this paper, we customize the cutting-edge Blended Latent Diffusion Model (BLDM) [21] to generate precisely aligned synthetic regions for industrial defect detection and localization.
+It is important to note that this work employs the diffusion model for generating synthetic images or image regions, just similar to the pioneering work [18, 47] proposed for the general editing tasks.
+Our method is irrelevant to the algorithms that use DPM to directly predict anomaly regions on the test images [48, 49, 50, 51].
 III The proposed method
@@
 I, a few defective (NG) samples $\mathcal{X}_{NG}=\{\mathbf{x}^{i}_{NG}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}\mid i=1,2,\dots,N_{NG}\}$ and the corresponding anomaly masks $\mathcal{M}_{NG}=\{\mathbf{M}^{i}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}\mid i=1,2,\dots,N_{NG}\}$ are given as the “seeds” of defect generation and usually $N_{NG}\ll N_{OK}$.
-The task of this paper is to generate the synthetic training set of defective samples as
+The goal of the proposed method is to generate a number of training samples with synthetic defective regions, i.e.,
 […]
-where $\mathcal{X}^{\ast}_{NG}=\{\mathbf{x}^{\ast i}_{NG}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}\mid i=1,2,\dots,N^{\ast}_{NG}\}$ and $\mathcal{M}^{\ast}_{NG}=\{\mathbf{M}^{\ast i}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}\mid i=1,2,\dots,N^{\ast}_{NG}\}$ comprise the obtained synthetic sample images and the corresponding label masks, respectively.
-Furthermore, the defect generation algorithm usually sets $N^{\ast}_{NG}\gg N_{NG}$ to ensure higher sample variations.
+where $\mathcal{X}^{\ast}_{NG}=\{\mathbf{x}^{\ast i}_{NG}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}\mid i=1,2,\dots,N^{\ast}_{NG}\}$ and $\mathcal{M}^{\ast}_{NG}=\{\mathbf{M}^{\ast i}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}\mid i=1,2,\dots,N^{\ast}_{NG}\}$ comprise the obtained synthetic samples and the corresponding defect masks, respectively.
+Furthermore, usually we set $N^{\ast}_{NG}\gg N_{NG}$ to ensure high pattern variation of the generated samples.
 In this paper, we propose to generate the synthetic samples using a diffusion model.
@@
 $\begin{split}&p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{t},t),\boldsymbol{\Sigma}_{\theta}(\mathbf{x}_{t},t))\\ &p_{\theta}(\mathbf{x}_{0:T})=p(\mathbf{x}_{T})\prod^{T}_{t=1}p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}),\end{split}$ (3)
-where $\boldsymbol{\mu}_{\theta}(\cdot)$ and $\boldsymbol{\Sigma}_{\theta}(\cdot)$ are the two unknown functions that are usually estimated as deep models.
+where $\boldsymbol{\mu}_{\theta}(\cdot)$ and $\boldsymbol{\Sigma}_{\theta}(\cdot)$ are the two unknown functions that are usually approximated as deep models.
 In the test phase, given different random Gaussian noise samples $\mathbf{x}^{i}_{T},i=1,\dots,N_{tst}$, one could obtain high quality and various samples $\mathbf{x}^{i}_{0},\forall i$ (usually RGB images) from the same domain of the training samples.
@@
 In this paper, we employ the Latent Diffusion Model (LDM) as the backbone of our method encouraged by its praiseworthy advantages [19].
-As the name indicating, the forward diffusion and reverse diffusion processes of LDM are all performed in the latent space.
-In general, given the input image $\mathbf{x}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$, its corresponding latent feature $\mathbf{z}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$ is obtained using the encoder function $\Omega(\cdot)$ as $\mathbf{z}=\Omega(\mathbf{x})$ while the image can also be recovered via the decoder function $\Phi(\cdot)$ as $\mathbf{x}=\Phi(\mathbf{z})$.
-Based on the latent feature $\mathbf{z}$ and some simplifications  [^1] Readers are referred to [16] and [33] for details from Eq.
-3, the reverse conditional probability in LDM is defined as
+As the name indicating, the forward and reverse diffusion processes of LDM are both performed in the lower-dimentional latent space.
+Given an input image $\mathbf{x}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$, its corresponding latent feature $\mathbf{z}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$ is obtained using the encoder function $\Omega(\cdot)$ as $\mathbf{z}=\Omega(\mathbf{x})$ while the image can also be recovered from $\mathbf{z}$ via the decoding operation as $\mathbf{x}=\Phi(\mathbf{z})$.
+Based on the latent feature $\mathbf{z}$ and some simplifications  [^1] Readers are referred to [19] and [38] for more details from Eq.
+3, the reverse conditional probability in latent space is revised as
 $\begin{split}&p_{\theta}(\mathbf{z}_{t-1}\mid\mathbf{z}_{t},t,\mathcal{C})=\\ &\mathcal{N}\left(\mathbf{z}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\right){\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathcal{C}),\sigma_{t}\text{I}\right),\end{split}$ (4)
@@
 III-B2 Linguistic prompts
-First, a linguistic prompt is obtained via changing the keywords of a following template sentence
+As an effective module, linguistic prompts are introduced to control and stabilize the generation process of diffusion models [19, 21].
+In this work, the linguistic prompt $y$ is obtained by specifying the keywords in the following template sentence
 […]
-where $obj$ and $def$ denote to the category of the current object (as the capsule in Fig.
+where the keywords $obj$ and $def$ denote the category of the current object (as the capsule in Fig.
 2) and defect type (as the black contamination on the capsule), respectively.
 One can see this template design is straightforward and thus the prompt is always easy to obtain.
-In this work, the sentence $y$ is then processed by the parameter-frozen text encoder $\tau(\cdot):\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}^{d_{lang}}$ which is proposed in the CLIP algorithm [48].
+The sentence $y$ is then processed by the parameter-frozen text encoder $\tau(\cdot):\mathbb{R}^{d_{y}}\rightarrow\mathbb{R}^{d_{lang}}$ of the CLIP model [52].
 In particular, as shown in Fig.
-2, the linguistic controlling information is fed into the generation model $\epsilon_{\theta}(\cdot)$ and the generation process is influenced by the K-Q-V attention operations [16].
-III-B3 Defect trimaps
-Additionally, recall that in practical anomaly detection tasks, the defect-free samples are usually sufficient in terms of both amount and variation, we thus propose to generate defective samples under the guidance of the randomly selected defect-free samples.
+2, the linguistic controlling information is fed into the generation model $\epsilon_{\theta}(\cdot)$ and the generation process is governed via a K-Q-V attention mechanism proposed in [19].
+III-B3 Defect trimap prompts
+Recall that in practical anomaly detection tasks, the defect-free samples are usually sufficient in terms of both amount and pattern variation, we thus propose to generate defective samples based on defect-free prototypes, rather than generating them from scratch.
 In a nutshell, a “trimap” $\Gamma\in\mathbb{R}^{H_{x}\times W_{x}}$ is designed to specify the desired location of the generated object and the defect, as shown in Fig.
 2.
-To obtain a proper trimap, we firstly estimate the foreground region i.e., the interested object, by using the method proposed in [11] and store the foreground mask into $\mathbf{F}\in\mathbb{B}^{H_{x}\times W_{x}}$ with $\mathbf{F}(x,y)=1$ indicating a foreground at the coordinate $[x,y]$.
-Secondly, a synthetic defect mask is generated based on the genuine anomaly mask set $\mathcal{M}_{NG}$ as
+To obtain a proper trimap, we firstly estimate the foreground region  [^2] The area corresponding to the interested object on a randomly selected defect-free image $\mathbf{x}_{OK}$ by using the method proposed in [14] and save the foreground mask as $\mathbf{F}\in\mathbb{B}^{H_{x}\times W_{x}}$ with $\mathbf{F}(x,y)=1$ indicating a foreground pixel at the coordinate $[x,y]$.
+Secondly, a synthetic defect mask is created based on a genuine anomaly mask randomly selected from set $\mathcal{M}_{NG}$.
+Formally, the generation of the novel anomaly mask writes:
 […]
@@
 The embedding feature is then fed into an encoder network $\hat{\epsilon}_{tri}(\cdot)$ that shares the same structure and initial parameters with the denoising encoder of $\epsilon_{\theta}(\cdot)$, except that the original “spatial attention” module is replaced with a self-attention process over the embedded trimap features.
-The deep features extracted from different layers of the denoising encoder are merged with the intermediate features of the denoising decoder of LDM via element-wise addition (see Fig.
+The deep features extracted from different layers of $\hat{\epsilon}_{tri}(\cdot)$ are then injected into the corresponding layers of the denoising decoder of $\epsilon_{\theta}(\cdot)$ via element-wise addition (see Fig.
 2).
-III-C Training Scheme
-Formally, our cross-modal controlling features is defined as $\mathcal{C}=\{\tau(y),\Gamma\}$ and the learning objective of LDM is defined as:
-$L_{LDM}\coloneqq\mathbb{E}_{\mathbf{z}_{t},\mathcal{C},\epsilon\thicksim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{C})\|^{2}_{2}\right].$ (8)
-In this work, the denoising model ${\epsilon}_{\theta}(\cdot)$, the embedding block $\zeta(\cdot)$ of trimap prompts and the encoder $\hat{\epsilon}_{tri}$ for the latent trimap feature are all learned via optimizing the loss function defined in Eq.
-8.
-However, different training schemes are applied for different modules.
-In specific, ${\epsilon}_{\theta}(\cdot)$ is firstly pre-trained by using the domain dataset (e.g.
-all the training images of all the subcategories of MVTec AD [19] for the experiments on this dataset) and then frozen during model learning for each specific subcategory; the latent feature encoder $\hat{\epsilon}_{tri}(\cdot)$ is initialized by the corresponding parameters of the original LDM model [16] and then fine-tuned for each subcategory based on the real defective samples; the convolutional block $\zeta(\cdot)$ is only learned based on the genuine anomaly samples with a random initialization.
-Apart from the reverse diffusion model, the text prompts $y$ are obtained by using the CLIP [48] and BLIP [49] algorithm for sufficiently various descriptions.
-The text encoder is copied from CLIP and frozen during training.
-As to the image encoder $\Omega(\cdot)$ and decoder $\Phi(\cdot)$, we follow [16] to inherit the deep models of the VQ-VAE algorithm [50] and fix them during the whole training stage.
+III-C Loss Function and Training Scheme
+Given the cross-modal controlling features $\{\tau(y),\zeta(\Gamma)\}$, the learning objective of the proposed model is defined as:
+$L_{LDM}\coloneqq\mathbb{E}_{\mathbf{z}_{t},\mathcal{C},\epsilon\thicksim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\tau(y),\hat{\epsilon}_{tri}(\zeta(\Gamma)))\|^{2}_{2}\right],$
+where ${\epsilon}_{\theta}(\cdot)$ denotes the denoising model and $\hat{\epsilon}_{tri}$ stands for the encoder for the latent trimap features.
+In this work, the involved modules are learned with different training schemes as described below.
+  • ${\epsilon}_{\theta}(\cdot)$ is pre-trained by using the domain dataset (e.g., all training images of all the subcategories of MVTec AD [22] for the experiments on this dataset).
+We keep ${\epsilon}_{\theta}(\cdot)$ frozen during model learning for each subcategory.
+  • The trimap feature encoder $\hat{\epsilon}_{tri}(\cdot)$ is initialized by the corresponding parameters of the original LDM model [19] and then fine-tuned for each subcategory based on the real defective samples
+  • The convolutional block $\zeta(\cdot)$ for trimap embedding is only learned based on the genuine anomaly samples from scratch.
+  • For the image encoder $\Omega(\cdot)$ and decoder $\Phi(\cdot)$, we follow [19] to deploy the corresponding modules of the VQ-VAE algorithm [53], which are fixed during the training.
+  • The text encoder $\tau(\cdot)$ is copied from CLIP and frozen during training.
 III-D Multi-Stage Denoising with Content Editing
-Inspired the seminal work, namely Blended Latent Diffusion (BLD) [17], we propose to perform feature editing in the inference phase.
-However, due to the task difference, we extensively modify the editing scheme of BLD for defect generation.
-First of all, given a defect-free image $\mathbf{x}_{OK}$ and a trimap $\Gamma$ defined before, the corresponding defect mask $\mathbf{M}^{\ast}_{NG}$ is firstly calculated as introduced in Sec.
-III-B.
+In this paper, we propose to perform feature editing in the inference phase by tailoring the editing scheme of Blended Latent Diffusion (BLD) [20] for defect generation.
+First of all, given a defect-free image $\mathbf{x}_{OK}$ and a trimap $\Gamma$, the corresponding defect mask $\mathbf{M}^{\ast}_{NG}$ is calculated as introduced in Sec.
+III-B3.
 Then the multi-stage denoising algorithm with content editing is performed and the workflow of this algorithm is illustrated in Alg.
-1 where $\odot$ indicates element-wise multiplication and $\neg$ stands for the element-wise “Not” operation on the binary map.
+1 where $\odot$ indicates element-wise multiplication and $\neg$ stands for the element-wise operation of logical “NOT” on the binary map.
+1
 Input: Defect-Free Image $\mathbf{x}_{OK}$, Trimap $\Gamma$, Defect Mask $\mathbf{M}^{\ast}_{NG}$, Language Prompt $y$, Decoder $\Psi(\cdot)$, Encoder $\Omega(\cdot)$, Range Numbers $T_{1}$, $T_{2}$ and $T_{3}$, Predefined Normal Distribution $\mathcal{N}(\mu_{z},\sigma^{2}_{z})$
+2
 Output: Blended Latent Feature $\mathbf{z}^{\ast}_{NG}$
-1 $\mathbf{z}_{OK}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}\leftarrow\Omega(\mathbf{x}_{OK})$
+3
+[…]
 5 $\mathbf{M}^{z}_{NG}\in\mathbb{B}^{H_{z}\times W_{z}}\leftarrow Dilate\&DownSample(\mathbf{M}^{\ast}_{NG})$
@@
 9 while $t>T_{2}+T_{3}$ do
+10
 11      $\mathbf{z}_{t-1}\leftarrow Denoise(\mathbf{z}_{t},\Gamma,y)$
@@
 13 end while
-10/* Latent Editing Stage */
+14/* Latent Editing Stage */
 15 while $t>T_{3}$ do
-12       $\mathbf{z}_{t}=\mathbf{z}_{t}\odot\mathbf{M}^{z}_{NG}+\mathbf{z}_{OK}\odot\neg\mathbf{M}^{z}_{NG}$ $\mathbf{z}_{t-1}\leftarrow Denoise(\mathbf{z}_{t},\Gamma,y)$
+16
+17      $\mathbf{z}_{t}=\mathbf{z}_{t}\odot\mathbf{M}^{z}_{NG}+\mathbf{z}_{OK}\odot\neg\mathbf{M}^{z}_{NG}$
+18      $\mathbf{z}_{t-1}\leftarrow Denoise(\mathbf{z}_{t},\Gamma,y)$
 19       $t=t-1$
 20 end while
-15/* Image Editing Stage */
+21/* Image Editing Stage */
 22 while $t>=0$ do
@@
 27 end while
-return $\mathbf{z}^{\ast}_{NG}=\mathbf{z}_{0}$
+28
+29return $\mathbf{z}^{\ast}_{NG}=\mathbf{z}_{0}$
 Algorithm 1 Multi-Stage Denoising with Editing
 As shown in Fig.
-2, the pipeline consistes of three stages in the denoising phase.
-The first stage is the conventional denosing process without content editing and takes $T_{1}$ steps.
-The second stage takes $T_{2}$ steps and the input feature $\mathbf{z}_{t}$ is merged with the feature of $\mathbf{x}_{OK}$, under the guidance of the dilated defect mask $\mathbf{M}^{\ast}_{NG}$.
+3, the pipeline consists of three stages in the denoising phase.
+The first stage takes $T_{1}$ steps to conduct the conventional denosing[38, 39, 19] without content editing.
+The second stage spends $T_{2}$ steps and the input feature $\mathbf{z}_{t}$ is merged with the feature of $\mathbf{x}_{OK}$, under the guidance of the dilated defect mask $\mathbf{M}^{\ast}_{NG}$.
 In the third stage, a novel content blending method proposed in this paper is conducted.
-In specific, the blending operation is performed in the pixel space rather than the latent space and the frozen encoder $\Omega(\cdot)$ and decoder $\Phi(\cdot)$ are used to switch the information between the two spaces.
+In specific, the blending operation is performed in the pixel space and the blended image is then remapped back to the latent space for denoising.
+The image encoder $\Omega(\cdot)$ and decoder $\Phi(\cdot)$ are used cooperatively to switch the information between the two spaces.
 Figure 3: The inference scheme of the proposed AdaBLDM algorithm.
@@
 III-E Online Decoder Adaptation
-Recall that the aim of our algorithm is to generate better synthetic defective images which are actually reconstructed by the decoder $\Phi(\cdot)$, we thus propose to fine tune the decoder for each generated sample to achieve higher image quality.
-Given the blended latent feature $\mathbf{z}^{\ast}_{NG}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$ obtained in Alg.
-1, the demanding defect mask $\mathbf{M}^{\ast}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}$ and the source image $\mathbf{x}_{OK}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$, one can fine tune the decoder model $\Phi(\cdot)$ by conducting the online update algorithm described in Alg.
+Considering that the image quality of the synthetic defective sample can be significantly influenced by the image decoder $\Phi(\cdot)$, we propose to fine-tune the decoder for each generated sample.
+Once the blended latent feature $\mathbf{z}^{\ast}_{NG}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$ is obtained from Alg.
+1, and given the defective mask $\mathbf{M}^{\ast}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}$ and the source image $\mathbf{x}_{OK}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$, we can fine-tune the decoder model $\Phi(\cdot)$ by conducting the online update algorithm as described in Alg.
+2, which shows that the online adaptation strategy encourages the decoder $\Phi(\cdot)$ generating defect-free pixels assembling to $\mathbf{x}_{OK}$ and simultaneously outputting the defective pixels guided by the original decoder.
+The conservative ratio $\lambda_{con}$ balances the two objectives.
+In the experiments of this paper, we demonstrate that this online-updated decoder can generate more realistic samples for AD training and thus leads to higher AD performance.
+1
+Input: Blended Latent Feature $\mathbf{z}^{\ast}_{NG}$, Defect-Free Image $\mathbf{x}_{OK}$, Maximum Step $T_{ft}$, Conservative Ratio $\lambda_{con}$, Anomaly Mask $\mathbf{M}^{\ast}_{NG}$, Decoder $\Phi(\cdot)$
 2
-Input: Blended Latent Feature $\mathbf{z}^{\ast}_{NG}$, Defect-Free Image $\mathbf{x}_{OK}$, Maximum Step $T_{ft}$, Conservative Ratio $\lambda_{con}$, Anomaly Mask $\mathbf{M}^{\ast}_{NG}$, Decoder $\Phi(\cdot)$
 Output: Refined Image $\mathbf{x}^{\ast}_{NG}$
-1 $\mathbf{x}^{\ast}_{NG}\leftarrow\Phi(\mathbf{z}^{\ast}_{NG}),~{}\bar{\mathbf{M}}^{\ast}_{NG}(i,j)\leftarrow\neg\mathbf{M}^{\ast}_{NG}(i,j)$
+3
+[…]
 5 $t=1$
 6 while $t<=T_{ft}$ do
+7
 8      $\tilde{\mathbf{x}}\leftarrow\Phi(\mathbf{z}^{\ast}_{NG})$
+9
 10      $L_{i}=\sum_{i,j}^{H_{x}W_{x}}\left(\tilde{\mathbf{x}}(i,j)-\mathbf{x}_{OK}(i,j)\right)^{2}\odot\bar{\mathbf{M}}^{\ast}_{NG}(i,j)$
+11
 12      $L_{d}=\sum_{i,j}^{H_{x}W_{x}}\left(\tilde{\mathbf{x}}(i,j)-\mathbf{x}^{\ast}_{NG}(i,j)\right)^{2}\odot\mathbf{M}^{\ast}_{NG}(i,j)$
+13
 14      $\Phi(\cdot)\leftarrow AdamW(L_{i}+\lambda_{con}L_{d},\Phi(\cdot))$
+15
 16      $t=t+1$
@@
 […]
-return $\mathbf{x}^{\ast}_{NG}$
+19
+20return $\mathbf{x}^{\ast}_{NG}$
 Algorithm 2 Online Decoder Adaptation
-One can see that the online adaptation strategy encourages the decoder $\Phi(\cdot)$ generating similar defect-free pixels as $\mathbf{x}_{OK}$ while outputting the defective pixels as the original decoder suggests.
-The conservative ratio $\lambda_{con}$ balance the two requirements.
-In the experiments of this work, we prove that the individually updated decoder can yield better samples and lead to higher AD performance.
 III-F Implementation Details
-In this paper, all the images are resize to $256\times 256$ following the setting of DeSTseg [6] which is a SOTA anomaly detection algorithm and adopted by this paper for evaluating the quality of generated samples.
+In this paper, all the images are resized to $256\times 256$ following the setting of DeSTseg [6] which is a SOTA anomaly detection algorithm adopted by this paper for evaluating the effectiveness of the generated samples.
 In this way, we set $H_{x}=W_{x}=256$ and $H_{z}=W_{z}=32$.
@@
 III-B3 is determined by a random rotation angle $\gamma\in[0^{\circ},360^{\circ}]$ and a random scaling factor $s\in[0.85,1.15]$.
-For the AD tasks on texture subcategories, such as a wooden or a fabric surface, all the pixels belong to the foreground region except the defective ones.
-We employ the AdamW optimizer [51] for updating the model parameters, with a default learning rate of $1\times 10^{-5}$, and default betas are set to $0.9$ and $0.999$, respectively.
-During training, there is a $10\%$ chance that empty text prompt is input to the model.
-During the inference stage, we adopt the DDIM [34] sampling method with the normal denoising step $T_{1}=50$, the latent blended step $T_{2}=30$, and pixel blended step $T_{3}=5$.
+For the AD tasks on texture subcategories, such as a wooden or a fabric surface, all the pixels are considered as foreground except the defective ones.
+We employed the AdamW optimizer [54] for updating the model parameters, with a default learning rate of $1\times 10^{-5}$, and default betas are set to $0.9$ and $0.999$, respectively.
+During training, there is a $10\%$ chance that an empty text prompt is input to the model.
+During the inference stage, we adopted the DDIM [39] sampling method with the normal denoising step $T_{1}=50$, the latent blended step $T_{2}=30$, and pixel blended step $T_{3}=5$.
 In the online decoder adaptation stage, $T_{ft}=200$ and the conservative ratio $\lambda_{con}=100$ to avoid overfitting to the anomaly-free region.
-The same optimizer as offline learning is used for this online adaptation except the learning rate is set to $1\times 10^{-4}$.
+The same optimizer as offline learning is used for this online adaptation except for the learning rate is set to $1\times 10^{-4}$.
 IV Experiments
@@
 The combination of the sophisticated (DeSTseg) and the naive (SVM) AD algorithms increases the comprehensiveness and thus the reliability of our experiment.
-IV-A2 Comparing Generation Methods
+IV-A2 Competitors
 We compare our defect generation algorithm with the cutting-edge generative models including DFM [17] and DCDGANc [16] which is built based on StarGAN [56] or StyleGANv2 [34] algorithms.
@@
 IV-A6 Training Settings for the Involved AD Models
-In this paper, we propose to evaluate the generated samples by using the simplest AD method and a sophisticated AD method. In this way, one can obtain a thorough examination of the generated samples.
-As one end of the spectrum, the SVM-based anomaly detection algorithm is extremely simple and thus can directly reflect the patch-wise quality of the generated images.
+In our experiments, the generated samples are evaluated by using the simplest AD method and a sophisticated AD method.
+As one end of the spectrum, the SVM-based anomaly detection algorithm is extremely simple and thus can directly reflect the primitive quality of the generated images, at the patch level.
 In specific, the SVM model based on the RBF kernel [64] is employed as the patch classifier.
@@
 As to the parameter setting, we by default set the $\gamma$ of the RBF kernel as $1\times 10^{-4}$ while the regularization parameter $C$ is set to $1$.
-As a SOTA anomaly detection method, DeSTseg [6] employs a discriminative model to detect anomalous regions based on comparing them and their reconstructed defect-free counterparts.
-Before training, we first generate $5000$ defective samples by using AdaBLDM and another $5000$ defective samples by using the data augmentation method of DeSTseg.
+Referring to DeSTseg [6] which employs a discriminative model to detect anomalous regions by comparing them with reconstructed defect-free counterparts, before training, we first generate $5000$ defective samples by using AdaBLDM and another $5000$ defective samples by using the data augmentation method of DeSTseg.
 We store the defective samples generated by AdaBLDM in $\mathcal{D}_{gen}$ and store those augmented by DeSTseg in $\mathcal{D}_{raw}$.
@@
 From top to bottom, the samples with genuine defects; the samples generated by using DCDGANc-StarGAN; the samples generated by using DCDGANc-StyleGAN; the samples generated with DFM and the samples generated by using our method.
-We further illustrate the comparison in a more fine-grained level in Fig.
+We further compare with existing methods in fine-grained details, and experimental results are reported in Fig.
 5.
-From the figure we can see that the proposed method is more well-controlled in generating the defect regions with the specific category and location.
+As shown in Fig.
+5, the proposed method is more well-controlled in generating the defect regions with the specific category and location.
+In addition, the generated images with synthetic defects on BTAD and KSDD2 are shown in Fig.
+6.
+It can be seen that besides MVTec AD, the image quality of AdaBLDM is maintained for the two different datasets.
 Figure 5: The fine-grained comparison in the generation quality on the Hazelnut subcategory of MVTec AD.
 Different kinds of defects are mimicked by the involved generation algorithms and their performances can be compared on a more detailed level.
-In addition, the generated images with synthetic defects on BTAD and KSDD2 are also shown in Fig.
-6.
-One can see that besides MVTec AD, the image quality of AdaBLDM is maintained for the two different datasets.
 Figure 6: The generation outputs of AdaBLDM on BTAD and KSDDV2 datasests.
@@
 7 shows the heat map of anomaly predicted by the DeSTseg algorithm.
-We can see that thanks to the higher quality of the simulated synthetics, the AdaBLDM-equipped AD algorithm can output the best heat maps for anomaly detection and localization.
+We can see that high-quality simulated synthetics benefits the AdaBLDM-equipped algorithm which can output competitive heat maps for anomaly detection and localization.
 Figure 7: Test results of different DeSTseg models learned based on the training sets obtained by using different generation methods.
@@
 IV-D Ablation Study
-To understand the success of the proposed AdaBLDM algorithm, we utilize ablation experiments to evaluate the specific contributions of the primary components which include:
+In this section, we systematically evaluate the contribution of each component in our method. The following experiments are conducted:
   • Pretrain with domain knowledge
@@
 V Conclusion & future work
-In this paper, we propose a sophisticated method to generate realistic defective samples for industrial anomaly detection and localization.
-Motivated by the great success of the diffusion model, we highly customize the Blended Latent Diffusion Model (BLDM) with several novel modifications for this specific task.
-The yielded algorithm is termed AdaBLDM and can generate better synthetic samples compared with the existing methods based on conventional “cut-and-paste” scheme or Generative Adversarial Networks (GANs).
-By applying the AdaBLDM for AD data augmentation, one can observe significant increase on the AD performances of SOTA algorithms.
-The AIGC algorithm proposed in this paper can be directly used as an effective data augmentation method in real-world applications and also sheds light on the influential factors for solving defect generalization task.
-In the future, one could obtain even better synthetic defective samples by introducing finer controlling signals such as the language prompts describing the details (shape, size etc.) about the demanding defects.
+In this paper, we propose an advanced method to generate realistic defective samples for industrial anomaly detection and localization.
+Inspired by the diffusion model, we customize the Blended Latent Diffusion Model (BLDM) to be well-suited to this specific task.
+The proposed algorithm, termed AdaBLDM, can generate better synthetic samples compared with existing methods based on conventional “cut-and-paste” schemes or Generative Adversarial Networks (GANs).
+By applying the AdaBLDM in anomaly detection data augmentation, we achieve a state-of-the-art performance of AD, thanks to the well-controlled defective regions aligned at the pixel level.
+In the future, we explore high-quality synthetic defective samples by introducing finer controlling signals such as the language prompts describing the details (shape, size, etc.) of the demanding defects.
 Another appealing direction of future work is to speed up the denoising process of the diffusion model so that more defective samples can be generated given a limited time budget, which is a crucial constraint in most practical scenarios.